# TalkingHeadBench: A Multi-Modal Benchmark & Analysis of Talking-Head DeepFake Detection

### v2

@@ 1 Introduction
Figure 1: Overview of the TalkingHeadBench dataset creation pipeline. We use portrait images from FFHQ [karras2019style] as source and audio or video from CelebV-HQ [zhu2022celebv] as driving signal to generate talking-head deepfake videos using 6 open-source generators for our core dataset. The images at the bottom show sample outputs from these six generators as well as from two additional emerging generators (MAGI-1 [magi1] and Hallo3 [cui2024hallo3]) used for evaluating generalization. We then perform multi-stage data curation to obtain 2994 high-quality deepfake videos along with 2312 real videos. Finally, we design three evaluation protocols to assess detector robustness under distribution shifts in identity and generator characteristics between training and testing sets.
@@
Deepfake technology, known for its ability to manipulate facial features and produce highly realistic synthetic videos, has seen rapid adoption in entertainment [liu_deepfacelab_2023, abdolahnejad_deep_2020], marketing [kaur_deepfake_2024], and personalized content creation [qi_mytimemachine_2024, qi_my3dgen_2025, wei_personalized_2025]. Yet, these advancements also introduce serious societal risks. Among deepfake types, talking-head deepfakes, which generate highly realistic head-and-shoulder videos of individuals speaking, are especially concerning. Driven by advanced generative models, these videos are potent tools for spreading misinformation, manipulating public discourse, and executing social engineering attacks [gambin_deepfakes_2024, saini_deepfake_2024, saif_fake_2024]. Their realism and ease of distribution make them particularly dangerous in sensitive domains like politics, finance, and media. In one real-world case, a finance worker was tricked into transferring \$25 million during a video call with a deepfake impersonating the company’s CFO [magramo_finance_2024]. These growing threats highlight the urgent need to understand talking-head deepfakes and to develop robust detection and mitigation strategies [gambin_deepfakes_2024, sirra_csslno_2024, liu_evolving_2025].
@@
In response to the growing threats posed by talking-head deepfakes, numerous facial deepfake benchmarks have been proposed (see Tab. 1), which played a central role in advancing detection methods. However, most existing benchmarks are not well-suited for the modern talking-head landscape. Most deepfake datasets primarily feature face-swapping techniques [rossler2019faceforensics++] rather than talking-head videos, which are generated using different techniques. Although large-scale and collected in the wild, these datasets often contain a high proportion of low-quality samples with noticeable artifacts [chandra2025deepfake, yan2024df40], making it easier to train detectors that ultimately perform poorly on more realistic deepfakes, which pose a greater threat. This leaves a gap in reliably measuring the performance of detectors against the more realistic manipulations produced by the recent talking-head generators. To address these gaps, we introduce TalkingHeadBench—the first manually curated, multi-generator, multi-modal benchmark designed specifically for evaluating detector robustness against state-of-the-art talking-head deepfakes.
@@
Our benchmark features high-quality synthetic videos generated with six academic talking-head generators using both audio- and video-driven diffusion and transformer-based models [drobyshev2024emoportraits, cui2024hallo2, xu2024hallo, wei2024aniportrait, guo2024liveportrait] in the training dataset. To further test generalization, we evaluate models against two emerging generators: an additional academic model (Hallo3 [cui2024hallo3]) and a commercial one (MAGI-1 [magi1]). To ensure dataset quality, we use a five-stage expert curation protocol that removes over 60% of generated videos based on pre-defined artifact criteria. The result is a challenging dataset of 2,994 high-quality talking-head deepfakes from eight academic and commercial generators. This emphasis on quality over quantity creates a more realistic evaluation setting. We also carefully design evaluation protocols to assess the generalization and robustness of state-of-the-art (SOTA) deepfake detectors [dong2023implicit, shao2025deepfake, xu2023tall, xu2024learning, liu2024lips] across identity and generator shifts in training and testing data. In addition to benchmarking, we present a detailed error analysis using Grad-CAM [Selvaraju_2019] to identify common failure cases of SOTA detectors on different generators and inform future improvements in detection strategies. We then note the limitations of SOTA detectors and identify areas of improvement, and provide guidelines on what performance future detectors should aspire to and how to analyze them.
@@
In summary, our benchmark demonstrates that while state-of-the-art detectors attain high accuracy on existing deepfake detection benchmarks [rossler2019faceforensics++], their performance drops markedly on TalkingHeadBench. This degradation stems from the inclusion of advanced multi-modal generators, deliberate shifts in identity and generator between training and testing splits, and careful curation that eliminates obvious artifacts in the deepfakes. Together, these factors highlight the need for developing more robust deepfake detectors and establish our benchmark as a foundation for driving future research in this direction.
@@
DeepFake-eval-2024 [chandra2025deepfake] N/A Video/Audio/Image 2025 1,208 767
FakeAVCeleb [khalid2021fakeavceleb] Multiple Audio/Video 2021 500 19,500
FaceForensics++ [rossler2019faceforensics++] Multiple Video 2019 1,000 4,000
LAV-DF [cai2022you] ✓* Single Audio/Video 2022 36,431 99,873
AV-Deepfake1M [cai2024av] ✓* Single Audio/Video 2023 286,721 860,039
PolyGlotFake [hou2024polyglotfakenovelmultilingualmultimodal] ✓* Multiple Audio 2024 766 14,472
DF-40  [yan2024df40] Multiple Video/Image 2024 1,590 1M+
@@ 2 Related Work
Deepfake generation has leveraged various facial manipulation techniques over the years. Among them, face swapping methods [nirkin2019fsgan, rosberg2023facedancer, chen2020simswap] are widely used, combining facial regions of various sources to create synthetic videos. They often introduce visible artifacts such as boundary inconsistency, lighting mismatch, or texture irregularity, which modern detectors can learn to exploit easily.
@@
Despite these advances, existing deepfake benchmarks remain limited in scope. As shown in Tab. 1, most prior datasets either focus on older GAN-based generators or target narrow manipulation types such as face swapping (e.g., FakeAVCeleb [khalid2021fakeavceleb] and FaceForensics++ [rossler2019faceforensics++]) or scrape social media content without controlling of generators (e.g., DeepFake-eval-2024 [chandra2025deepfake]). While newer datasets like LAV-DF [cai2023glitch], AV-Deepfake1M [cai2024av], and PolyGlobFake [hou2024polyglotfakenovelmultilingualmultimodal] incorporate talking-head deepfakes, they mostly emphasize lip-sync-driven synthesis and lack broader modality or generator diversity. Further, these datasets are often massive but uncurated, resulting in many low-quality samples. In contrast, our dataset is built using a curated collection of cutting-edge diffusion-based talking-head generators spanning both audio and video-driven modalities. Through manual inspection, we discarded over 60% of generated videos due to visible artifacts (see Tab. 4), ensuring that only high-quality, realistic samples remain—thereby offering a more meaningful and challenging benchmark.
Recent detection methods (see Tab. 3), though effective on static and unimodal datasets like FaceForensics++ [rossler2019faceforensics++], have not been thoroughly evaluated against multi-modal diffusion-based deepfakes. To fill this gap, we introduce a comprehensive benchmark that reflects the current generation landscape. Our results show that SOTA detectors, which achieve near-perfect accuracy on traditional benchmarks like FaceForensics++, struggle significantly when tested on our dataset, highlighting poor robustness to identity and generator shifts and calling for stronger generalization capabilities.
@@
Image, Audio Diffusion Hallo [xu2024hallo] arXiv24
Image, Audio Diffusion Hallo2 [cui2024hallo2] ICLR25
Image, Audio, Text Diffusion Hallo3 [cui2024hallo3] CVPR25
Image, Video Diffusion X-Portrait [xie2024x] SIGGRAPH24
Image, Video Non-diffusion LivePortrait [guo2024liveportrait] arXiv24
Image, Video Non-diffusion EMOPortraits [drobyshev2024emoportraits] CVPR24
Image, Video Non-diffusion MCNet [hong2023implicit] ICCV23
Image, Video Non-diffusion FSRT [rochow2024fsrt] CVPR24
Image, Audio/Video Diffusion AniPortrait [wei2024aniportrait] arXiv24
Image, Audio/Video Diffusion SkyReels-A1 [qiu2025skyreelsa1expressiveportraitanimation] arXiv25
Image, Text Diffusion MAGI-1 [magi1] N/A
@@
Image CNN SBI [shiohara2022detecting] CVPR22
Image Neural Network CADDM [dong2023implicit] CVPR23
Image Neural Network HiFiNet[hifi_net_xiaoguo]//HiFiNet++[xiao_hifinet_plusplus] CVPR23/IJCV24
Image Transformer DeepFake-Adapter [shao2025deepfake] IJCV25
Image Transformer GenConVit [deressa2025genconvitdeepfakevideodetection] arXiv25
Image Transformer RFFR [shi2025real] PR25
Image Neural network LAA-Net [nguyen2024laa] CVPR24
Video Transformer TALL [xu2023tall]/TALL++ [xu2024learning] ICCV23/IJCV24
Video CNN AltFreezing [wang2023altfreezing] CVPR23
Video CNN/Transformer NACO [zhang2024learning] ECCV24
Video Neural network MM-Det [song2025learningmultimodalforgeryrepresentation] NeurIPS24
Video Neural network StyleFlow [choi2024exploiting] CVPR24
Audio/Video Transformer LipFD [liu2024lips] NeurIPS 2024
@@ 3.1 Source Data and Generation Overview
To create a diverse deepfake benchmark dataset, we utilize publicly available datasets containing images from FFHQ (Flickr-Faces-HQ) [karras2019style] and videos of real people and audio of people talking from CelebV-HQ [zhu2022celebv], processed by a diverse suite of contemporary deepfake generators. We select six open-source generators: Hallo [xu2024hallo], Hallo2 [cui2024hallo2], AniPortrait (Audio-driven) (AniAudio) [wei2024aniportrait], AniPortrait (Video-driven) (AniVideo) [wei2024aniportrait], LivePortrait [guo2024liveportrait], and EMOPortraits [drobyshev2024emoportraits], aiming for diversity in both diffusion-based and non-diffusion-based frameworks and driving signal modalities in both video and audio, as summarized in Tab. 2. We choose these generators because they are the latest SoTA generators with publicly available code covering varying modalities and frameworks.
Additionally, we test generalization against two entirely unseen models: Hallo3 [cui2024hallo3], a recent academic generator, and MAGI-1 [magi1], a commercial text-to-video model. These are used exclusively in our test set to evaluate out-of-distribution robustness. Owing to computational and cost limitations, these test sets are relatively small in size, but they nevertheless provide a meaningful probe of model generalization. We also explore other generators such as X-Portrait [xie2024x] and MCNet [hong2023implicit], but do not include them in our dataset due to significant artifacts observed.
@@
Data Type # Total # Train # Test Ratio Removed Hallo [xu2024hallo] 420 303 117 $\sim 42\%$ Hallo2 [cui2024hallo2] 432 327 105 $\sim 60\%$ AniPortrait (Audio) [wei2024aniportrait] 542 396 146 $\sim 66.7\%$ AniPortrait (Video) [wei2024aniportrait] 422 314 108 $\sim 80\%$ LivePortrait [guo2024liveportrait] 529 352 177 $\sim 60\%$ EMOPortraits [drobyshev2024emoportraits] 573 381 192 $\sim 50\%$ MAGI-1 [magi1] 66 0 66 $\sim 18\%$ Hallo3 [cui2024hallo3] 10 0 10 $\sim 10\%$ TOTAL Fakes 2994 2073 921 $\sim 63\%$
@@ 4.1 Evaluation Setup
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.90 0.55 0.25 0.17 0.83 0.39 0.39 0.21 0.94 0.65 0.64 0.16 0.96 0.45 0.45 0.11 0.94 0.27 0.23 0.17 0.95 0.59 0.42 0.13 HiFi-Net [hifi_net_xiaoguo] 0.95 0.47 0.27 0.09 0.95 0.66 0.41 0.09 0.93 0.39 0.19 0.11 0.92 0.22 0.12 0.12 0.81 0.03 0.01 0.25 0.90 0.26 0.04 0.12 AltFreezing [wang2023altfreezing] 1.00 0.93 0.71 0.03 1.00 0.91 0.75 0.05 1.00 0.97 0.97 0.03 1.00 0.95 0.88 0.05 1.00 0.93 0.89 0.04 1.00 0.94 0.89 0.04 TALL [xu2023tall] 1.00 1.00 1.00 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 1.00 0.99 0.98 0.01 1.00 0.98 0.98 0.05 1.00 0.97 0.96 0.01 LipFD [liu2024lips] 0.98 0.75 0.58 0.05 0.99 0.74 0.38 0.05 0.98 0.72 0.54 0.05 0.99 0.83 0.62 0.04 0.99 0.78 0.50 0.05 0.98 0.84 0.56 0.05 DF-Adapter [shao2025deepfake] 0.99 0.58 0.06 0.01 1.00 1.00 0.15 0.01 1.00 0.99 0.11 0.01 1.00 0.99 0.98 0.01 1.00 0.99 0.25 0.01 1.00 0.98 0.41 0.01 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.97 0.65 0.26 0.14 0.97 0.55 0.19 0.15 0.97 0.63 0.26 0.14 0.96 0.57 0.21 0.15 0.93 0.34 0.11 0.14 0.91 0.28 0.04 0.15
@@
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.88 0.32 0.02 0.18 0.94 0.27 0.18 0.09 0.97 0.61 0.52 0.08 0.94 0.37 0.14 0.12 0.79 0.02 0.00 0.21 0.98 0.71 0.63 0.06 HiFi-Net [hifi_net_xiaoguo] 0.88 0.08 0.01 0.16 0.82 0.16 0.06 0.18 0.82 0.02 0.00 0.20 0.48 0.02 0.00 0.45 0.58 0.01 0.00 0.34 0.85 0.01 0.01 0.20 AltFreezing [wang2023altfreezing] 1.00 0.97 0.82 0.02 1.00 0.94 0.66 0.02 1.00 0.98 0.87 0.01 0.98 0.71 0.53 0.06 0.96 0.69 0.52 0.08 1.00 0.94 0.78 0.02 TALL [xu2023tall] 1.00 1.00 0.98 0.01 1.00 0.99 0.98 0.02 1.00 1.00 1.00 0.00 0.97 0.89 0.81 0.08 0.97 0.70 0.40 0.15 1.00 1.00 0.98 0.01 LipFD [liu2024lips] 0.99 0.76 0.51 0.05 0.97 0.62 0.30 0.09 0.98 0.70 0.47 0.08 0.87 0.50 0.32 0.17 0.89 0.30 0.08 0.23 0.99 0.81 0.63 0.05 DF-Adapter [shao2025deepfake] 1.00 1.00 0.83 0.01 1.00 1.00 1.00 0.00 1.00 1.00 0.73 0.01 0.97 0.80 0.68 0.15 0.99 0.93 0.80 0.13 1.00 1.00 0.99 0.01 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.93 0.39 0.04 0.12 0.95 0.36 0.14 0.11 0.94 0.34 0.08 0.15 0.97 0.51 0.25 0.06 0.89 0.15 0.07 0.06 0.88 0.23 0.11 0.16
@@
AniAudio [wei2024aniportrait] AniVideo [wei2024aniportrait] Hallo [xu2024hallo] Hallo2 [cui2024hallo2] EMOPortraits [drobyshev2024emoportraits] LivePortrait [guo2024liveportrait] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [dong2023implicit] 0.82 0.25 0.21 0.023 0.98 0.60 0.60 0.06 0.95 0.55 0.20 0.11 0.93 0.17 0.17 0.10 0.75 0.03 0.03 0.24 0.95 0.52 0.41 0.09 HiFi-Net [hifi_net_xiaoguo] 0.88 0.03 0.01 0.17 0.86 0.27 0.11 0.16 0.84 0.00 0.00 0.18 0.48 0.00 0.00 0.44 0.61 0.01 0.00 0.34 0.84 0.01 0.00 0.27 AltFreezing [wang2023altfreezing] 1.00 0.87 0.78 0.03 1.00 0.94 0.88 0.03 1.00 0.95 0.89 0.02 0.99 0.85 0.64 0.04 0.95 0.58 0.26 0.08 0.99 0.90 0.56 0.03 TALL [xu2023tall] 1.00 1.00 0.99 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 0.97 0.84 0.82 0.06 0.97 0.48 0.48 0.15 1.00 0.93 0.92 0.02 LipFD [liu2024lips] 0.99 0.72 0.62 0.04 0.98 0.71 0.56 0.08 0.98 0.72 0.34 0.08 0.89 0.57 0.40 0.15 0.91 0.30 0.14 0.22 0.98 0.77 0.61 0.06 DF-Adapter [shao2025deepfake] 1.00 0.93 0.16 0.02 1.00 0.99 0.94 0.01 1.00 1.00 0.95 0.00 0.98 0.88 0.80 0.09 0.98 0.70 0.53 0.14 1.00 0.95 0.56 0.03 MM-Det [song2025learningmultimodalforgeryrepresentation] 0.94 0.49 0.06 0.13 0.96 0.40 0.16 0.12 0.93 0.33 0.06 0.18 0.98 0.55 0.27 0.06 0.87 0.06 0.03 0.18 0.85 0.08 0.03 0.17
@@
(i) How does the performance of the SOTA detectors change from FaceForensics++ [rossler2019faceforensics++], where detectors have near-perfect accuracy, to TalkingHeadBench?

### v3

@@ 1 Introduction
Figure 1: Overview of the TalkingHeadBench dataset creation pipeline. We use portrait images from FFHQ [18] as source and audio or video from CelebV-HQ [44] as driving signal to generate talking-head deepfake videos using 6 open-source generators for our core dataset. The images at the bottom show sample outputs from these six generators as well as from two additional emerging generators (MAGI-1 [30] and Hallo3 [8]) used for evaluating generalization. We then perform multi-stage data curation to obtain 2994 high-quality deepfake videos along with 2312 real videos. Finally, we design three evaluation protocols to assess detector robustness under distribution shifts in identity and generator characteristics between training and testing sets.
@@
Deepfake technology, known for its ability to manipulate facial features and produce highly realistic synthetic videos, has seen rapid adoption in entertainment [liu_deepfacelab_2023, abdolahnejad_deep_2020], marketing [kaur_deepfake_2024], and personalized content creation [24, 25, wei_personalized_2025]. Yet, these advancements also introduce serious societal risks. Among deepfake types, talking-head deepfakes, which generate highly realistic head-and-shoulder videos of individuals speaking, are especially concerning. Driven by advanced generative models, these videos are potent tools for spreading misinformation, manipulating public discourse, and executing social engineering attacks [gambin_deepfakes_2024, saini_deepfake_2024, saif_fake_2024]. Their realism and ease of distribution make them particularly dangerous in sensitive domains like politics, finance, and media. In one real-world case, a finance worker was tricked into transferring \$25 million during a video call with a deepfake impersonating the company’s CFO [magramo_finance_2024]. These growing threats highlight the urgent need to understand talking-head deepfakes and to develop robust detection and mitigation strategies [gambin_deepfakes_2024, sirra_csslno_2024, liu_evolving_2025].
@@
In response to the growing threats posed by talking-head deepfakes, numerous facial deepfake benchmarks have been proposed (see Tab. 1), which played a central role in advancing detection methods. However, most existing benchmarks are not well-suited for the modern talking-head landscape. Most deepfake datasets primarily feature face-swapping techniques [29] rather than talking-head videos, which are generated using different techniques. Although large-scale and collected in the wild, these datasets often contain a high proportion of low-quality samples with noticeable artifacts [4, 42], making it easier to train detectors that ultimately perform poorly on more realistic deepfakes, which pose a greater threat. This leaves a gap in reliably measuring the performance of detectors against the more realistic manipulations produced by the recent talking-head generators. To address these gaps, we introduce TalkingHeadBench—the first manually curated, multi-generator, multi-modal benchmark designed specifically for evaluating detector robustness against state-of-the-art talking-head deepfakes.
@@
Our benchmark features high-quality synthetic videos generated with six academic talking-head generators using both audio- and video-driven diffusion and transformer-based models [11, 7, 39, 37, 12] in the training dataset. To further test generalization, we evaluate models against two emerging generators: an additional academic model (Hallo3 [8]) and a commercial one (MAGI-1 [30]). To ensure dataset quality, we use a five-stage expert curation protocol that removes over 60% of generated videos based on pre-defined artifact criteria. The result is a challenging dataset of 2,994 high-quality talking-head deepfakes from eight academic and commercial generators. This emphasis on quality over quantity creates a more realistic evaluation setting. We also carefully design evaluation protocols to assess the generalization and robustness of state-of-the-art (SOTA) deepfake detectors [10, 32, 40, 41, 20] across identity and generator shifts in training and testing data. In addition to benchmarking, we present a detailed error analysis using Grad-CAM [31] to identify common failure cases of SOTA detectors on different generators and inform future improvements in detection strategies. We then note the limitations of SOTA detectors and identify areas of improvement, and provide guidelines on what performance future detectors should aspire to and how to analyze them.
@@
In summary, our benchmark demonstrates that while state-of-the-art detectors attain high accuracy on existing deepfake detection benchmarks [29], their performance drops markedly on TalkingHeadBench. This degradation stems from the inclusion of advanced multi-modal generators, deliberate shifts in identity and generator between training and testing splits, and careful curation that eliminates obvious artifacts in the deepfakes. Together, these factors highlight the need for developing more robust deepfake detectors and establish our benchmark as a foundation for driving future research in this direction.
@@
DeepFake-eval-2024 [4] N/A Video/Audio/Image 2025 1,208 767
FakeAVCeleb [19] Multiple Audio/Video 2021 500 19,500
FaceForensics++ [29] Multiple Video 2019 1,000 4,000
LAV-DF [3] ✓* Single Audio/Video 2022 36,431 99,873
AV-Deepfake1M [1] ✓* Single Audio/Video 2023 286,721 860,039
PolyGlotFake [16] ✓* Multiple Audio 2024 766 14,472
DF-40  [42] Multiple Video/Image 2024 1,590 1M+
@@ 2 Related Work
Deepfake generation has leveraged various facial manipulation techniques over the years. Among them, face swapping methods [23, 28, 5] are widely used, combining facial regions of various sources to create synthetic videos. They often introduce visible artifacts such as boundary inconsistency, lighting mismatch, or texture irregularity, which modern detectors can learn to exploit easily.
@@
Despite these advances, existing deepfake benchmarks remain limited in scope. As shown in Tab. 1, most prior datasets either focus on older GAN-based generators or target narrow manipulation types such as face swapping (e.g., FakeAVCeleb [19] and FaceForensics++ [29]) or scrape social media content without controlling of generators (e.g., DeepFake-eval-2024 [4]). While newer datasets like LAV-DF [2], AV-Deepfake1M [1], and PolyGlobFake [16] incorporate talking-head deepfakes, they mostly emphasize lip-sync-driven synthesis and lack broader modality or generator diversity. Further, these datasets are often massive but uncurated, resulting in many low-quality samples. In contrast, our dataset is built using a curated collection of cutting-edge diffusion-based talking-head generators spanning both audio and video-driven modalities. Through manual inspection, we discarded over 60% of generated videos due to visible artifacts (see Tab. 4), ensuring that only high-quality, realistic samples remain—thereby offering a more meaningful and challenging benchmark.
Recent detection methods (see Tab. 3), though effective on static and unimodal datasets like FaceForensics++ [29], have not been thoroughly evaluated against multi-modal diffusion-based deepfakes. To fill this gap, we introduce a comprehensive benchmark that reflects the current generation landscape. Our results show that SOTA detectors, which achieve near-perfect accuracy on traditional benchmarks like FaceForensics++, struggle significantly when tested on our dataset, highlighting poor robustness to identity and generator shifts and calling for stronger generalization capabilities.
@@
Image, Audio Diffusion Hallo [39] arXiv24
Image, Audio Diffusion Hallo2 [7] ICLR25
Image, Audio, Text Diffusion Hallo3 [8] CVPR25
Image, Video Diffusion X-Portrait [38] SIGGRAPH24
Image, Video Non-diffusion LivePortrait [12] arXiv24
Image, Video Non-diffusion EMOPortraits [11] CVPR24
Image, Video Non-diffusion MCNet [15] ICCV23
Image, Video Non-diffusion FSRT [27] CVPR24
Image, Audio/Video Diffusion AniPortrait [37] arXiv24
Image, Audio/Video Diffusion SkyReels-A1 [26] arXiv25
Image, Text Diffusion MAGI-1 [30] N/A
@@
Image CNN SBI [34] CVPR22
Image Neural Network CADDM [10] CVPR23
Image Neural Network HiFiNet[14]//HiFiNet++[13] CVPR23/IJCV24
Image Transformer DeepFake-Adapter [32] IJCV25
Image Transformer GenConVit [9] arXiv25
Image Transformer RFFR [33] PR25
Image Neural network LAA-Net [22] CVPR24
Video Transformer TALL [40]/TALL++ [41] ICCV23/IJCV24
Video CNN AltFreezing [36] CVPR23
Video CNN/Transformer NACO [43] ECCV24
Video Neural network MM-Det [35] NeurIPS24
Video Neural network StyleFlow [6] CVPR24
Audio/Video Transformer LipFD [20] NeurIPS 2024
@@ 3.1 Source Data and Generation Overview
To create a diverse deepfake benchmark dataset, we utilize publicly available datasets containing images from FFHQ (Flickr-Faces-HQ) [18] and videos of real people and audio of people talking from CelebV-HQ [44], processed by a diverse suite of contemporary deepfake generators. We select six open-source generators: Hallo [39], Hallo2 [7], AniPortrait (Audio-driven) (AniAudio) [37], AniPortrait (Video-driven) (AniVideo) [37], LivePortrait [12], and EMOPortraits [11], aiming for diversity in both diffusion-based and non-diffusion-based frameworks and driving signal modalities in both video and audio, as summarized in Tab. 2. We choose these generators because they are the latest SoTA generators with publicly available code covering varying modalities and frameworks.
Additionally, we test generalization against two entirely unseen models: Hallo3 [8], a recent academic generator, and MAGI-1 [30], a commercial text-to-video model. These are used exclusively in our test set to evaluate out-of-distribution robustness. Owing to computational and cost limitations, these test sets are relatively small in size, but they nevertheless provide a meaningful probe of model generalization. We also explore other generators such as X-Portrait [38] and MCNet [15], but do not include them in our dataset due to significant artifacts observed.
@@
Data Type # Total # Train # Test Ratio Removed Hallo [39] 420 303 117 $\sim 42\%$ Hallo2 [7] 432 327 105 $\sim 60\%$ AniPortrait (Audio) [37] 542 396 146 $\sim 66.7\%$ AniPortrait (Video) [37] 422 314 108 $\sim 80\%$ LivePortrait [12] 529 352 177 $\sim 60\%$ EMOPortraits [11] 573 381 192 $\sim 50\%$ MAGI-1 [30] 66 0 66 $\sim 18\%$ Hallo3 [8] 10 0 10 $\sim 10\%$ TOTAL Fakes 2994 2073 921 $\sim 63\%$
@@ 4.1 Evaluation Setup
AniAudio [37] AniVideo [37] Hallo [39] Hallo2 [7] EMOPortraits [11] LivePortrait [12] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [10] 0.90 0.55 0.25 0.17 0.83 0.39 0.39 0.21 0.94 0.65 0.64 0.16 0.96 0.45 0.45 0.11 0.94 0.27 0.23 0.17 0.95 0.59 0.42 0.13 HiFi-Net [14] 0.95 0.47 0.27 0.09 0.95 0.66 0.41 0.09 0.93 0.39 0.19 0.11 0.92 0.22 0.12 0.12 0.81 0.03 0.01 0.25 0.90 0.26 0.04 0.12 AltFreezing [36] 1.00 0.93 0.71 0.03 1.00 0.91 0.75 0.05 1.00 0.97 0.97 0.03 1.00 0.95 0.88 0.05 1.00 0.93 0.89 0.04 1.00 0.94 0.89 0.04 TALL [40] 1.00 1.00 1.00 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 1.00 0.99 0.98 0.01 1.00 0.98 0.98 0.05 1.00 0.97 0.96 0.01 LipFD [20] 0.98 0.75 0.58 0.05 0.99 0.74 0.38 0.05 0.98 0.72 0.54 0.05 0.99 0.83 0.62 0.04 0.99 0.78 0.50 0.05 0.98 0.84 0.56 0.05 DF-Adapter [32] 0.99 0.58 0.06 0.01 1.00 1.00 0.15 0.01 1.00 0.99 0.11 0.01 1.00 0.99 0.98 0.01 1.00 0.99 0.25 0.01 1.00 0.98 0.41 0.01 MM-Det [35] 0.97 0.65 0.26 0.14 0.97 0.55 0.19 0.15 0.97 0.63 0.26 0.14 0.96 0.57 0.21 0.15 0.93 0.34 0.11 0.14 0.91 0.28 0.04 0.15
@@
AniAudio [37] AniVideo [37] Hallo [39] Hallo2 [7] EMOPortraits [11] LivePortrait [12] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [10] 0.88 0.32 0.02 0.18 0.94 0.27 0.18 0.09 0.97 0.61 0.52 0.08 0.94 0.37 0.14 0.12 0.79 0.02 0.00 0.21 0.98 0.71 0.63 0.06 HiFi-Net [14] 0.88 0.08 0.01 0.16 0.82 0.16 0.06 0.18 0.82 0.02 0.00 0.20 0.48 0.02 0.00 0.45 0.58 0.01 0.00 0.34 0.85 0.01 0.01 0.20 AltFreezing [36] 1.00 0.97 0.82 0.02 1.00 0.94 0.66 0.02 1.00 0.98 0.87 0.01 0.98 0.71 0.53 0.06 0.96 0.69 0.52 0.08 1.00 0.94 0.78 0.02 TALL [40] 1.00 1.00 0.98 0.01 1.00 0.99 0.98 0.02 1.00 1.00 1.00 0.00 0.97 0.89 0.81 0.08 0.97 0.70 0.40 0.15 1.00 1.00 0.98 0.01 LipFD [20] 0.99 0.76 0.51 0.05 0.97 0.62 0.30 0.09 0.98 0.70 0.47 0.08 0.87 0.50 0.32 0.17 0.89 0.30 0.08 0.23 0.99 0.81 0.63 0.05 DF-Adapter [32] 1.00 1.00 0.83 0.01 1.00 1.00 1.00 0.00 1.00 1.00 0.73 0.01 0.97 0.80 0.68 0.15 0.99 0.93 0.80 0.13 1.00 1.00 0.99 0.01 MM-Det [35] 0.93 0.39 0.04 0.12 0.95 0.36 0.14 0.11 0.94 0.34 0.08 0.15 0.97 0.51 0.25 0.06 0.89 0.15 0.07 0.06 0.88 0.23 0.11 0.16
@@
AniAudio [37] AniVideo [37] Hallo [39] Hallo2 [7] EMOPortraits [11] LivePortrait [12] Detector AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier AUC T1 T0.1 Brier CADDM [10] 0.82 0.25 0.21 0.023 0.98 0.60 0.60 0.06 0.95 0.55 0.20 0.11 0.93 0.17 0.17 0.10 0.75 0.03 0.03 0.24 0.95 0.52 0.41 0.09 HiFi-Net [14] 0.88 0.03 0.01 0.17 0.86 0.27 0.11 0.16 0.84 0.00 0.00 0.18 0.48 0.00 0.00 0.44 0.61 0.01 0.00 0.34 0.84 0.01 0.00 0.27 AltFreezing [36] 1.00 0.87 0.78 0.03 1.00 0.94 0.88 0.03 1.00 0.95 0.89 0.02 0.99 0.85 0.64 0.04 0.95 0.58 0.26 0.08 0.99 0.90 0.56 0.03 TALL [40] 1.00 1.00 0.99 0.01 1.00 1.00 1.00 0.01 1.00 0.99 0.99 0.01 0.97 0.84 0.82 0.06 0.97 0.48 0.48 0.15 1.00 0.93 0.92 0.02 LipFD [20] 0.99 0.72 0.62 0.04 0.98 0.71 0.56 0.08 0.98 0.72 0.34 0.08 0.89 0.57 0.40 0.15 0.91 0.30 0.14 0.22 0.98 0.77 0.61 0.06 DF-Adapter [32] 1.00 0.93 0.16 0.02 1.00 0.99 0.94 0.01 1.00 1.00 0.95 0.00 0.98 0.88 0.80 0.09 0.98 0.70 0.53 0.14 1.00 0.95 0.56 0.03 MM-Det [35] 0.94 0.49 0.06 0.13 0.96 0.40 0.16 0.12 0.93 0.33 0.06 0.18 0.98 0.55 0.27 0.06 0.87 0.06 0.03 0.18 0.85 0.08 0.03 0.17
@@
(i) How does the performance of the SOTA detectors change from FaceForensics++ [29], where detectors have near-perfect accuracy, to TalkingHeadBench?